# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v2

@@ Appendix E Additional Results
$\sum_{i=1}^{N}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i}\Big{)}\cdot h({\bm{x}},{\bm{\eta}}_{i})$
$\hskip 85.35826pt=\sum_{i=1}^{N_{*}}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}+c^{*}_{i}\Big{)}\cdot h({\bm{x}},{\bm{\eta}}^{*}_{i}).$ (72)
Note that since the expert function $h(\cdot,{\bm{\eta}})$ satisfies the conditions in Definition 3.2, then given an arbitrary $N^{\prime}\in\mathbb{N}$, then the set $\{h({\bm{x}},{\bm{\eta}}^{\prime}_{i}):i\in[N^{\prime}]\}$, where ${\bm{\eta}}^{\prime}_{1},\ldots,{\bm{\eta}}^{\prime}_{N^{\prime}}$ are distinct vectors, is linearly independent for almost every ${\bm{x}}$. If $N\neq N_{*}$, then there exists some $i\in[N]$ such that ${\bm{\eta}}_{i}\neq{\bm{\eta}}^{*}_{j}$ for any $j\in[N_{*}]$. This implies that $\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i}\Big{)}=0$, which is a contradiction. Thus, we must have that $N=N_{*}$. As a result, we get that
$\Big{\{}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i}\Big{)}:i\in[N]\Big{\}}=\Big{\{}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}+c^{*}_{i}\Big{)}:i\in[N_{*}]\Big{\}},$
@@
for almost every ${\bm{x}}$ for any $i\in[N_{*}]$. It is worth noting that the $\mathrm{Softmax}$ function is invariant to translations, then equation (73) indicates that ${\bm{A}}_{i}={\bm{A}}^{*}_{i}+\mathbf{T}_{2}$ ${\bm{b}}_{i}={\bm{b}}^{*}_{i}+\mathbf{t}_{1}$ and $c_{i}=c^{*}_{i}+t_{0}$ for some $\mathbf{T}_{2}\in\mathbb{R}^{d\times d}$, $\mathbf{t}_{1}\in\mathbb{R}^{d}$ and $t_{0}\in\mathbb{R}$. However, from the assumptions ${\bm{A}}_{k}={\bm{A}}^{*}_{k}$, ${\bm{b}}_{k}={\bm{b}}^{*}_{k}=\mathbf{0}_{d}$ and $c_{k}=c^{*}_{k}=0$, we deduce that $\mathbf{T}_{2}=\mathbf{0}_{d\times d}$, $t_{1}=\mathbf{0}_{d}$ and $t_{0}=0$. Consequently, we get that ${\bm{A}}_{i}={\bm{A}}^{*}_{i}$, ${\bm{b}}_{i}={\bm{b}}^{*}_{i}$ and $c_{i}=c^{*}_{i}$ for any $i\in[N_{*}]$. Then, equation (E) can be rewritten as
$\sum_{i=1}^{N_{*}}\exp(c_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}\Big{)}h({\bm{x}},{\bm{\eta}}_{i})=\sum_{i=1}^{N_{*}}\exp(c^{*}_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}\Big{)}h({\bm{x}},{\bm{\eta}}^{*}_{i}),$ (74)
for almost every ${\bm{x}}$. Next, we denote $P_{1},P_{2},\ldots,P_{M}$ as a partition of the index set $[N_{*}]$, where $M\leq N_{*}$, such that $\exp(c_{i})=\exp(c^{*}_{i^{\prime}})$ for any $i,i^{\prime}\in P_{j}$ and $j\in[N_{*}]$. On the other hand, when $i$ and $i^{\prime}$ do not belong to the same set $P_{j}$, we let $\exp(c_{i})\neq\exp(c_{i^{\prime}})$. Thus, we can reformulate equation (74) as
$\sum_{j=1}^{M}\sum_{i\in{P}_{j}}\exp(c_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}\Big{)}h({\bm{x}},{\bm{\eta}}_{i})$
$\hskip 85.35826pt=\sum_{j=1}^{M}\sum_{i\in{P}_{j}}\exp(c^{*}_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}\Big{)}h({\bm{x}},{\bm{\eta}}^{*}_{i}),$
for almost every ${\bm{x}}$. Recall that ${\bm{A}}_{i}={\bm{A}}^{*}_{i}$, ${\bm{b}}_{i}={\bm{b}}^{*}_{i}$ and $c_{i}=c^{*}_{i}$ for any $i\in[N_{*}]$, then the above equation implies that
@@
for almost every ${\bm{x}}$ for any $j\in[m]$. As a consequence,
@@
The proof of Proposition E.2 can be done in a similar fashion to that of Proposition E.1.
Appendix F Experimental Details
F.1 Verification of Theoretical Results.
F.2 Performance of Active-Attention Mechanism.
F.2.1 Dataset Details
CIFAR-10 Dataset. CIFAR-10 [18] is an established computer-vision dataset used for object recognition. It consists of 60,000 32$\times$32 color images containing one of 10 object classes ("plane", "car", "bird", "cat", "deer", "dog", "frog", "horse", "ship", "truck"), with 6000 images per class.
ImageNet Dataset. We use the ImageNet database from ILSVRC2012 [36] that contains $1.28M$ training images and $50K$ validation images, where the task is to classify images into 1,000 distinct categories, using a vast dataset of over 1.2 million training images and 150,000 validation and test images sourced from the ImageNet database.
@@ F.3 Model Details
Language Models  We used the small versions of language models developed by [37]. For both the Transformer and Performer models, the dimensions of the key, value, and query were set to 128, and the context length for training and evaluation was configured to 256. Each model was assigned 8 self-attention heads, with the feedforward network (FFN) having a hidden dimension of 2048. The number of attention layers was set to 16. These models are used for the WikiText-103 experiment depicted in Figure 2 (c).
Time Series Models  We use the supervised PatchTST model with the default parameter configurations proposed in [28]. The look-back window is set to 336, with a patch length of 16 and a stride of 8. For the Illness dataset, the prediction length is 36, while for the rest of the datasets, it is set to 192. For the Transformer model, we use a standard self-attention transformer with an encoder-decoder architecture, consisting of 2 encoder layers and 1 decoder layer. The model uses 8 attention heads and a feed-forward network (FFN) hidden dimension of 2048.
F.4 Quadratic Gating vs. Linear Gating.
F.4.1 Dataset details.
FineWeb-Edu Dataset. FineWeb-Edu 10BT [30] is a large-scale dataset with 10 billion tokens curated from diverse educational sources like textbooks and academic publications, designed to enhance language models for educational tasks. Its high-quality filtering and deduplication processes make it ideal for training relatively small-scale language models.
F.4.2 Model details.
GPT2-MoE. We examined the MoE adaptation of the GPT-2 [33] transformer structure. Specifically, we substituted the FFN layers with an MoE layer consisting of 8 experts. To ensure a fair comparison between MoE models and the dense model, we set the hidden size of the FFN layer ($d_{\mathrm{ff}}$) so that the active parameters during inference are roughly equivalent. We also set the quadratic gate rank to 32 to limit the number of additional gating parameters. Refer to Table 5 for comprehensive details on the parameter counts. Details on the hyperparameters related to the architecture and training process are provided in Table 6.
@@
Appendix G Computational Overhead of Quadratic Gating
@@
In this section, we analyze the computational and memory overhead introduced by incorporating quadratic gating. Consider an MoE layer with $N$ two-layer MLP experts with hidden size of $d_{\mathrm{ff}}$. Note that each expert has $2d\times d_{\mathrm{ff}}$ parameters. The linear gating network has $d$ parameters per expert while the introduction of quadratic terms in the gating network increases the parameter count to at most $d+d(d+1)/2$ per expert. Therefore, in each MoE layer the ratio of additional gating parameters to the total number of parameters is $d/4d_{\mathrm{ff}}$.
@@
Parameter count overhead. Note that this parameter count overhead is not necessarily negligible; however, this can be effectively managed by using low-rank embeddings for second-order terms in the router. For example, in Mixtral 8x7B, $d_{\mathrm{ff}}=3.5\times d$ which results in approximately $7\%$ increase in MoE layer parameter count if we use quadratic gating. The total parameter count overhead is almost $2.1$B, which is about $4\%$ of total number of parameters ($47$B) or $16\%$ of total number of active parameters ($13$B). In particular, assuming that the query and key linear embeddings in the Att-MoE framework are low-rank with rank $r\ll d$, the number of additional parameters per expert can be reduced to $(\frac{N+1}{N})(r\times d)$. Therefore, the ratio of additional gating parameters to the total number of parameters in the MoE layers can be as low as $(\frac{N+1}{N})(r/2d_{\mathrm{ff}})$. For the Mixtral 8x7B example, assuming $r=128$, the total parameter count overhead reduces to 150M parameters, which is only $0.3\%$ of all parameters and $1.1\%$ of the active parameters, making it relatively insignificant. Similarly, in our experiments with GPT2 level models, we used $r=32$, which leads to roughly $2.3$M additional total parameters which is $1.4\%$ of the total parameters and $1.8\%$ of active parameters.
@@
Computational overhead. The number of FLOPs is usually considered proportional to the number of non-embedding parameters. In the case of the Mixtral 8x7B example, utilizing low-rank embeddings with $r=128$ for both query and key vectors within the Att-MoE framework adds 150M parameters, resulting in just a $1\%$ increase in the number of FLOPs.
@@
Memory overhead. Since all parameters must be loaded into memory for inference or training, the memory usage of MoE models is proportional to their total number of parameters. For example, in the Mixtral 8x7B model, using low-rank embeddings with $r=128$ in the Att-MoE framework, which leads to merely a $0.3\%$ increase in memory overhead.

### v3

@@ Appendix B Identifiability of Quadratic Gating MoE
$\sum_{i=1}^{N}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i}\Big{)}\cdot\mathcal{E}({\bm{x}},{\bm{\eta}}_{i})$
$\hskip 85.35826pt=\sum_{i=1}^{N^{*}}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}+c^{*}_{i}\Big{)}\cdot\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{i}).$ (71)
Note that since the expert function $h(\cdot,{\bm{\eta}})$ satisfies the conditions in Definition 3.2, then given an arbitrary $N^{\prime}\in\mathbb{N}$, then the set $\{\mathcal{E}({\bm{x}},{\bm{\eta}}^{\prime}_{i}):i\in[N^{\prime}]\}$, where ${\bm{\eta}}^{\prime}_{1},\ldots,{\bm{\eta}}^{\prime}_{N^{\prime}}$ are distinct vectors, is linearly independent for almost every ${\bm{x}}$. If $N\neq N^{*}$, then there exists some $i\in[N]$ such that ${\bm{\eta}}_{i}\neq{\bm{\eta}}^{*}_{j}$ for any $j\in[N^{*}]$. This implies that $\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i}\Big{)}=0$, which is a contradiction. Thus, we must have that $N=N^{*}$. As a result, we get that
$\Big{\{}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i}\Big{)}$ $:i\in[N]\Big{\}}$
=$=\Big{\{}\mathrm{Softmax}\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}+c^{*}_{i}\Big{)}:i\in[N^{*}]\Big{\}},$
@@
for almost every ${\bm{x}}$ for any $i\in[N^{*}]$. It is worth noting that the $\mathrm{Softmax}$ function is invariant to translations, then equation (72) indicates that ${\bm{A}}_{i}={\bm{A}}^{*}_{i}+\mathbf{T}_{2}$ ${\bm{b}}_{i}={\bm{b}}^{*}_{i}+\mathbf{t}_{1}$ and $c_{i}=c^{*}_{i}+t_{0}$ for some $\mathbf{T}_{2}\in\mathbb{R}^{d\times d}$, $\mathbf{t}_{1}\in\mathbb{R}^{d}$ and $t_{0}\in\mathbb{R}$. However, from the assumptions ${\bm{A}}_{k}={\bm{A}}^{*}_{k}$, ${\bm{b}}_{k}={\bm{b}}^{*}_{k}=\mathbf{0}_{d}$ and $c_{k}=c^{*}_{k}=0$, we deduce that $\mathbf{T}_{2}=\mathbf{0}_{d\times d}$, $t_{1}=\mathbf{0}_{d}$ and $t_{0}=0$. Consequently, we get that ${\bm{A}}_{i}={\bm{A}}^{*}_{i}$, ${\bm{b}}_{i}={\bm{b}}^{*}_{i}$ and $c_{i}=c^{*}_{i}$ for any $i\in[N^{*}]$. Then, equation (B) can be rewritten as
$\sum_{i=1}^{N^{*}}\exp(c_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}\Big{)}\mathcal{E}({\bm{x}},{\bm{\eta}}_{i})=\sum_{i=1}^{N^{*}}\exp(c^{*}_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}\Big{)}\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{i}),$ (73)
for almost every ${\bm{x}}$. Next, we denote $P_{1},P_{2},\ldots,P_{M}$ as a partition of the index set $[N^{*}]$, where $M\leq N^{*}$, such that $\exp(c_{i})=\exp(c^{*}_{i^{\prime}})$ for any $i,i^{\prime}\in P_{j}$ and $j\in[N^{*}]$. On the other hand, when $i$ and $i^{\prime}$ do not belong to the same set $P_{j}$, we let $\exp(c_{i})\neq\exp(c_{i^{\prime}})$. Thus, we can reformulate equation (73) as
$\sum_{j=1}^{M}\sum_{i\in{P}_{j}}\exp(c_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}\Big{)}\mathcal{E}({\bm{x}},{\bm{\eta}}_{i})$
$\hskip 85.35826pt=\sum_{j=1}^{M}\sum_{i\in{P}_{j}}\exp(c^{*}_{i})\exp\Big{(}{\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}\Big{)}\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{i}),$
for almost every ${\bm{x}}$. Recall that ${\bm{A}}_{i}={\bm{A}}^{*}_{i}$, ${\bm{b}}_{i}={\bm{b}}^{*}_{i}$ and $c_{i}=c^{*}_{i}$ for any $i\in[N^{*}]$, then the above equation implies that
@@
for almost every ${\bm{x}}$ for any $j\in[m]$.
As a consequence,
@@
The proof of Proposition B.2 can be done in a similar fashion to that of Proposition B.1.
Appendix C Experimental Details
C.1 Performance of Active-Attention Mechanism
C.1.1 Dataset Details
CIFAR-10 Dataset. CIFAR-10 krizhevsky2009learning is an established computer-vision dataset used for object recognition. It consists of 60,000 32$\times$32 color images containing one of 10 object classes ("plane", "car", "bird", "cat", "deer", "dog", "frog", "horse", "ship", "truck"), with 6000 images per class.
ImageNet Dataset. We use the ImageNet database from ILSVRC2012 (russakovsky2015imagenet, ) that contains $1.28M$ training images and $50K$ validation images, where the task is to classify images into 1,000 distinct categories, using a vast dataset of over 1.2 million training images and 150,000 validation and test images sourced from the ImageNet database.
@@ C.1.2 Model Details
Language Models  We used the small versions of language models developed by (schlag2021linear, ). For both the Transformer and Performer models, the dimensions of the key, value, and query were set to 128, and the context length for training and evaluation was configured to 256. Each model was assigned 8 self-attention heads, with the feedforward network (FFN) having a hidden dimension of 2048. The number of attention layers was set to 16. These models are used for the WikiText-103 experiment depicted in Figure 2 (c).
Time Series Models  We use the supervised PatchTST model with the default parameter configurations proposed in Yuqietal-2023-PatchTST . The look-back window is set to 336, with a patch length of 16 and a stride of 8. For the Illness dataset, the prediction length is 36, while for the rest of the datasets, it is set to 192. For the Transformer model, we use a standard self-attention transformer with an encoder-decoder architecture, consisting of 2 encoder layers and 1 decoder layer. The model uses 8 attention heads and a feed-forward network (FFN) hidden dimension of 2048.
C.2 Quadratic Gating vs. Linear Gating
C.2.1 Dataset details
FineWeb-Edu Dataset. FineWeb-Edu 10BT (penedo2024fineweb, ) is a large-scale dataset with 10 billion tokens curated from diverse educational sources like textbooks and academic publications, designed to enhance language models for educational tasks. Its high-quality filtering and deduplication processes make it ideal for training relatively small-scale language models.
C.2.2 Model details.
GPT2-MoE. We examined the MoE adaptation of the GPT-2 (radford2019language, ) transformer structure. Specifically, we substituted the FFN layers with an MoE layer consisting of 8 experts. To ensure a fair comparison between MoE models and the dense model, we set the hidden size of the FFN layer ($d_{\mathrm{ff}}$) so that the active parameters during inference are roughly equivalent. We also set the quadratic gate rank to 32 to limit the number of additional gating parameters. Refer to Table 6 for comprehensive details on the parameter counts. Details on the hyperparameters related to the architecture and training process are provided in Table 7.
@@ C.2.1 Dataset details
Appendix D Computational Overheads of Quadratic Gating
@@
In this section, we analyze the computational and memory overheads introduced by incorporating quadratic gating. Consider an MoE layer with $N$ two-layer MLP experts with hidden size of $d_{\mathrm{ff}}$. Note that each expert has $2d\times d_{\mathrm{ff}}$ parameters. The linear gating network has $d$ parameters per expert while the introduction of quadratic terms in the gating network increases the parameter count to at most $d+d(d+1)/2$ per expert. Therefore, in each MoE layer the ratio of additional gating parameters to the total number of parameters is $d/4d_{\mathrm{ff}}$.
@@
Parameter count overheads. Note that this parameter count overheads is not necessarily negligible; however, this can be effectively managed by using low-rank embeddings for second-order terms in the router. For example, in Mixtral 8x7B, $d_{\mathrm{ff}}=3.5\times d$ which results in approximately $7\%$ increase in MoE layer parameter count if we use quadratic gating. The total parameter count overheads is almost $2.1$B, which is about $4\%$ of total number of parameters ($47$B) or $16\%$ of total number of active parameters ($13$B). In particular, assuming that the query and key linear embeddings in the Att-MoE framework are low-rank with rank $r\ll d$, the number of additional parameters per expert can be reduced to $(\frac{N+1}{N})(r\times d)$. Therefore, the ratio of additional gating parameters to the total number of parameters in the MoE layers can be as low as $(\frac{N+1}{N})(r/2d_{\mathrm{ff}})$. For the Mixtral 8x7B example, assuming $r=128$, the total parameter count overheads reduces to 150M parameters, which is only $0.3\%$ of all parameters and $1.1\%$ of the active parameters, making it relatively insignificant. Similarly, in our experiments with GPT2 level models, we used $r=32$, which leads to roughly $2.3$M additional total parameters which is $1.4\%$ of the total parameters and $1.8\%$ of active parameters.
@@
Computational overheads. The number of FLOPs is usually considered proportional to the number of non-embedding parameters. In the case of the Mixtral 8x7B example, utilizing low-rank embeddings with $r=128$ for both query and key vectors within the Att-MoE framework adds 150M parameters, resulting in just a $1\%$ increase in the number of FLOPs.
@@
Memory overheads. Since all parameters must be loaded into memory for inference or training, the memory usage of MoE models is proportional to their total number of parameters. For example, in the Mixtral 8x7B model, using low-rank embeddings with $r=128$ in the Att-MoE framework, which leads to merely a $0.3\%$ increase in memory overheads.
Appendix E Ablation Study on the Number of Attention Heads
We present results from an ablation study examining the effect of the number of attention heads. We select the best-performing activation function for each dataset, along with the linear activation, and evaluate their performance across varying numbers of attention heads using the best base model for that dataset. The results in Table 8 show that performance improves as the number of heads increases in most situations. This is particularly evident in image-based tasks and most time-series forecasting tasks, supporting the intuition that increased model complexity can enhance performance, especially for large-scale datasets. Information on model architectures, hyperparameters, and training details of experiments on non-linear gating functions can be found in Tables 9 - 12.
Table 8: Performance across varying numbers of attention heads.
Dataset Activation # of Attention Heads
2 4 8 12 16
CIFAR-10 Linear 79.41 82.68 84.98 85.22 85.88
GeLU 82.04 83.28 85.79 86.50 87.26
ImageNet Linear 73.34 79.35 82.31 82.44 84.42
GeLU 78.55 80.79 84.47 85.36 86.13
Weather Linear 0.198 0.197 0.197 0.196 0.195
Tanh 0.187 0.187 0.186 0.187 0.184
Traffic Linear 0.385 0.383 0.384 0.382 0.382
Tanh 0.379 0.375 0.375 0.373 0.372
Electricity Linear 0.153 0.152 0.152 0.153 0.151
Tanh 0.142 0.141 0.141 0.140 0.141
Illness Linear 1.517 1.474 1.462 1.458 1.449
Tanh 1.498 1.447 1.453 1.442 1.439
ETTh1 Linear 0.419 0.414 0.414 0.413 0.412
Tanh 0.417 0.410 0.412 0.414 0.408
ETTh2 Linear 0.342 0.338 0.339 0.335 0.333
Sigmoid 0.329 0.325 0.323 0.318 0.317
ETTm1 Linear 0.328 0.331 0.327 0.327 0.325
Tanh 0.322 0.325 0.324 0.325 0.326
ETTm2 Linear 0.212 0.212 0.212 0.211 0.212
Tanh 0.212 0.212 0.212 0.211 0.212
Table 9: Summary of ViT-Small Model for CIFAR-10
Model Architecture
Base Model Vision Transformer (ViT)
Patch Embedding Shifted Patch Tokenization (SPT)
Transformer Blocks 6
Attention Mechanism Local Self-Attention (LSA)
FFN Activation GELU
Classifier Head LayerNorm + Linear Layer
Hyperparameters
Input Image Size 32x32
Patch Size 4x4
Embedding Dimension 512
Attention Heads 8
Dimension per Head 64
Transformer MLP Dimension 512
LSA Activation Function GELU
Pooling CLS Token
Dropout 0.1
Embedding Dropout 0.1
Training Details
Dataset CIFAR-10
Epochs 400
Batch Size 512
Optimizer Adam
Learning Rate 1e-4
LR Scheduler Cosine Annealing
Loss Function Cross-Entropy Loss
Data Augmentation RandAugment, RandomCrop, RandomHorizontalFlip
Mixed Precision Enabled
Table 10: Summary of ViT Model for ImageNet
Model Architecture
Base Model Vision Transformer (ViT)
Patch Embedding Linear Projection
Transformer Blocks 12
Attention Mechanism Multi-Head Self-Attention
Attention Activation Configurable (e.g., GELU, ReLU, etc.)
FFN Activation GELU
Classifier Head LayerNorm + Linear Layer
Hyperparameters
Input Image Size 224x224
Patch Size 16x16
Embedding Dimension 768
Attention Heads 12
Dimension per Head 64
Transformer MLP Dimension 3072 (4 * Embedding Dim)
Pooling CLS Token
Dropout 0.1
Embedding Dropout 0.1
Training Details
Dataset ImageNet
Epochs 300
Batch Size 8
Optimizer Adam
Learning Rate 1e-4
LR Scheduler Cosine Annealing
Loss Function Cross-Entropy Loss
Data Augmentation RandAugment, RandomResizedCrop, RandomHorizontalFlip
Mixed Precision Enabled
Table 11: Summary of PatchTST model architecture, hyperparameters, and training details.
Model Architecture
Component Description
Input Processing The input time series is first normalized using Reversible Instance Normalization (RevIN).
Patching The normalized series is partitioned into overlapping patches.
Patch Embedding Each patch is linearly projected to an embedding.
Positional Encoding Learned positional encodings are added to the patch embeddings.
Encoder A standard Transformer encoder architecture is used to process the sequence of patch embeddings.
Attention Standard multi-head self-attention. The implementation includes residual attention connections as in Realformer. A non-standard application of a non-linear activation function is applied to the value projection (V) inside the attention mechanism.
Feed-Forward Network Each encoder layer contains a position-wise feed-forward network.
Head The output embeddings from the encoder are flattened and passed through a linear layer to produce a forecast. An option for channel-independent heads is available.
Hyperparameters
Hyperparameter Value
Sequence Length 336
Prediction Length 192
Patch Length 16
Stride 8
Encoder Layers 3
Attention Heads A range is tested, e.g., 2, 4, 8, 16
Model Dimension 16
FFN Dimension 128
Dropout 0.3
FC Dropout 0.3
Head Dropout 0.0
FFN Activation GeLU
Attention Value Activation Configurable (e.g., GELU, ReLU, etc.)
Table 12: PatchTST Training Details
Detail Configuration
Optimizer Adam
Learning Rate 0.0001
Learning Rate Scheduler OneCycleLR
Loss Function Mean Squared Error (MSE)
Batch Size 16
Epochs 100
Early Stopping Enabled with a patience of 100 epochs.